# Also at ]IFISC (UIB-CSIC), Instituto de Fisica Interdisciplinar y Sistemas Complejos - Palma de Mallorca, Spain

### v1

@@ Abstract
The Mpemba effect refers to the surprising observation where, under certain conditions, a far-from-equilibrium state can relax toward equilibrium faster than a state closer to equilibrium. A paradigmatic example is provided by the curious fact that hot water can sometimes freeze faster than cold water. The Mpemba effect has intrigued scientists since long time and has been predicted and observed in a variety of classical and quantum systems. Recently, the search for Mpemba-like effects of purely quantum nature has raised a major interest. Here we predict the emergence of Mpemba effect in the quantum optics context exploiting non-classical states of light. By analyzing the decay dynamics of photon fields in a leaky optical resonator or waveguide, it is demonstrated that the Mpemba effect emerges in the context of the quantum nature of light. Specifically, the relaxation dynamics is strongly influenced by the photon statistics of the initially trapped light field. The Mpemba effect is observed when comparing the decay dynamics of classical light fields (coherent states) with certain non-classical states, such as squeezed states and Schrödinger cat states.
@@
The Mpemba effect (ME) S1 ; S2 ; S2b ; S3 describes the counterintuitive scenario where, under certain conditions, a far-from-equilibrium state can relax toward equilibrium faster than a state closer to equilibrium. A paradigmatic example, known since long time, is provided by the curious fact that warmer water can sometimes freeze faster than colder water S1 ; S2 ; S2b . The ME has sparked an intense debate for decades S3 ; S3b and, in spite of being extensively studied in many classical systems S2b ; S3 ; S3b ; S4 ; S5 ; S6 ; S3c ; S6b ; S6c , it remains a fascinating and somewhat enigmatic phenomenon that continues to intrigue scientists. Recently, the ME is being intensely explored in the microscopic world S7 ; S8 ; S9 ; S9b ; S10 ; S11 ; S11b ; S12 ; S12b ; S13 ; S14 ; S14b ; S15 ; S16 ; S17 ; fuffa ; Busch ; noo ; S17b , with the prediction and observation of quantum versions of the ME. For example, in the dynamics of closed many-body systems an anomalous relaxation process can occur in which greater initial symmetry breaking of a subsystem leads to faster symmetry restoration S11 ; S12 ; S12b ; S14b . Such results are stimulating the search for Mpemba-like phenomena of purely quantum origin, which could be of relevance in quantum thermodynamics and could inspire innovative technologies in areas ranging from computing to energy storage S17b . While the ME has been demonstrated in various physical systems, it has largely been overlooked by the photonics community. Recently, the ME has been predicted to arise in the diffusion process of light waves in certain photonic lattices under incoherent dynamics S18 . Such previous study indicated that, rather surprisingly, tightly-localized light waves can diffuse and reach uniform distribution in the lattice faster than weakly-confined fields, thus providing the first example of ME involving light waves. However, this effect concerning light diffusion is purely classical, prompting the key and intriguing question: do Mpemba-like effects for quantum light exist, i.e. originating from the quantum nature of light, which is fundamental to modern integrated photonic quantum systems?
In this Letter we predict the emergence of the Mpemba effect in quantum optics context originating from the quantum nature of light book , which does not have any classical counterpart. Specifically, we consider the decay dynamics of photon fields from a leaky optical resonator or waveguide R1 ; R2 ; R3 ; R3b ; R4 ; R5 ; Haus1 ; R6 ; R7 ; R8 (Fig.1), and show that the relaxation dynamics is strongly affected by the photon statistics of initially trapped light field. In particular, it is shown that ME arising from the quantum nature of light is observed when comparing the decay dynamics of coherent states (classical states of light) with certain non-classical states of light, such as Fock states, squeezed states and Schrödinger cat states.
@@
At the classical level the initial state is determined solely by the initial electric field amplitude $A(t=0)$ of the resonator mode, and the decay dynamics is the same regardless of the initial field amplitude. To unveil the occurrence of ME in the photon decay dynamics, we should consider the quantum nature of light. To this aim, let us assume that at initial time the photon field is described by the factorized state $|\psi(t=0)\rangle=|\psi_{S}(0)\rangle_{S}\otimes|0\rangle_{B}$, where $|0\rangle_{B}$ is the vacuum state of the reservoir (radiation) modes and
@@
and $A=A(t)$ is the classical decay law, i.e. the solution to Eq.(4) with the initial condition $A(t=0)=1$. The derivation of Eq.(9) is given in the Appendix A.
Since $A(t)\rightarrow 0$ as $t\rightarrow\infty$, for any arbitrary initial probability amplitudes $\{p_{n}\}$ one has $\lim_{t\rightarrow\infty}\hat{\rho}_{S}(t)=|0\rangle_{S}\langle 0|\equiv\hat{\rho}_{S}^{\infty}$, i.e. as expected the asymptotic reduced density operator converges toward the vacuum state. The condition $A(t)\neq 0$ at any finite time $t$ ensures that in the transient relaxation process toward the stationary state it is not possible for the system to transiently cross the stationary state, which would introduce some ambiguities in the definition of the ME. The main point is that the relaxation dynamics strongly depends on the initial probability amplitudes $\{p_{n}\}$ of the photon field, which could result in the ME under suitable choice of initial states. To quantify how much the quantum state $\hat{\rho}_{S}(t)$ is far from $\hat{\rho}_{S}^{\infty}$ during the relaxation process, a distance measure $D(t)$ for density operators in Hilbert space has to be defined. Such a definition is not unique and ambiguities in the onset of ME could arise S5 ; S6c , however it is generally argued that any distance measure satisfying some minimal properties should result in a unique definition of the ME S5 . For quantum systems, the trace distance $D_{T}(t)$ and the Hilbert-Schmidt distance $D_{HS}(t)$ are generally used as possible metrics of distances in Hilbert space S9 ; S9b ; N1 ; N2 . They are defined by the relations
@@
Figure 2: Mpemba effect in photon decay dynamics with non-classical light. The figure compares the relaxation dynamics in the resonator/waveguide for an initial single-photon Fock state $|1\rangle_{S}=\hat{a}^{{\dagger}}|0\rangle_{S}$ (state I) and for the coherent state $|\alpha_{0}^{(coh)}\rangle_{S}$ (state II) with amplitude $\alpha_{0}=1$. Panels (a,b) show the behaviors of trace distance ($D_{T}$) and Hilbert-Schmidt distance ($D_{HS}$) versus normalized time $t/\tau_{c}$. A classical exponential decay law for $|A(t)|^{2}$ has been assumed, with photon lifetime $\tau_{c}$. The crossing between the two curves I and II, which arises at about $t/\tau_{c}\simeq 0.34$ for both metrics, is the signature of the ME. Panel (c) shows a few snapshots of the reduced density matrix elements $|\rho^{(S)}_{n,m}(t)|$ at some increasing times [from left to right: $t/\tau_{c}=0$ (initial states), $t/\tau_{c}=0.6$, $t/\tau_{c}=1.6$ and $t/\tau_{c}=3$]. Panels in the upper row refer to state I (Fock state decay), whereas panels in the lower row refer to state II (coherent state decay).
@@
Figure 3: Same as Fig.2, except that state I is the Schrödinger cat state $|\psi_{S}(t=0)\rangle_{S}=\mathcal{N}(|\alpha_{0}^{(coh)}\rangle_{S}+|-\alpha_{0}^{(coh)}\rangle_{S})$ with $\alpha_{0}=1.5$ ($\mathcal{N}$ is a normalization constant). As compared to Fig.2, the crossing time of the two curves I and II depends on the metrics ($t/\tau_{c}\simeq 1.7$ in (a) and $t/\tau_{c}\simeq 1$ in (b)).
Figure 4: Same as Fig.2, except that state II is the squeezed vacuum state $|\psi_{S}(t=0)\rangle_{S}=\exp[(1/2)(\xi^{*}\hat{a}^{2}-\xi\hat{a}^{{\dagger}2})]|0\rangle_{S}$ with squeezing parameter $\xi=1$.
@@
In conclusion, we predicted the existence of a quantum Mpemba effect for light by considering photon decay dynamics of non-classical states of light. The analysis shows that certain non-classical states of light, such as photon number (Fock) states, decay faster than classical (coherent) light states when the full degrees of freedom of the photon field (i.e. reduced density matrix) are monitored. Our results unveil a type of photonic Mpemba effect of purely quantum nature, extending such a counterintuitive and curious phenomenon to the quantum optics context with potential interest in advanced quantum photonic technologies, such as quantum computing, quantum simulations, quantum metrology and quantum communications fuffa ; S17b , where fast relaxation dynamics could be desirable.

### v2

@@ Abstract
The Mpemba effect refers to the surprising observation where, under certain conditions, a far-from-equilibrium state can relax toward equilibrium faster than a state closer to equilibrium. A paradigmatic example is provided by the curious fact that hot water can sometimes freeze faster than cold water. The Mpemba effect has intrigued scientists since long time and has been predicted and observed in a variety of classical and quantum systems. Recently, the search for Mpemba-like effects of purely quantum nature has raised a major interest. Here we predict the emergence of Mpemba effect in the quantum optics context exploiting non-classical states of light. By analyzing the decay dynamics of photon fields in a leaky optical resonator or waveguide, it is demonstrated that bosonic Mpemba effect emerges in the context of the quantum nature of light. Specifically, the relaxation dynamics is strongly influenced by the photon statistics of the initially trapped light field. The Mpemba effect is observed when comparing the decay dynamics of classical light fields (coherent states) with certain non-classical states, such as Fock states, squeezed states and Schrödinger cat states.
@@
The Mpemba effect (ME) S1 ; S2 ; S2b ; S3 describes the counterintuitive scenario where, under certain conditions, a far-from-equilibrium state can relax toward equilibrium faster than a state closer to equilibrium. A paradigmatic example, known since long time, is provided by the curious fact that warmer water can sometimes freeze faster than colder water S1 ; S2 ; S2b . The ME has sparked an intense debate for decades S3 ; S3b and, in spite of being extensively studied in many classical systems S2b ; S3 ; S3b ; S4 ; S5 ; S6 ; S3c ; S6b ; S6c , it remains a fascinating and somewhat enigmatic phenomenon that continues to intrigue scientists. Recently, the ME is being intensely explored in the microscopic world S7 ; S8 ; S9 ; S9b ; S10 ; S11 ; S11b ; S12 ; S12b ; S13 ; S14 ; S14b ; S15 ; S16 ; S17 ; fuffa ; Busch ; noo ; cin1 ; cin2 ; S17b , with the prediction and observation of quantum versions of the ME. For example, in the dynamics of closed many-body systems an anomalous relaxation process can occur in which greater initial symmetry breaking of a subsystem leads to faster symmetry restoration S11 ; S12 ; S12b ; S14b . Such results are stimulating the search for Mpemba-like phenomena of purely quantum origin, which could be of relevance in quantum thermodynamics and could inspire innovative technologies in areas ranging from computing to energy storage S17b . While the ME has been demonstrated in various physical systems, it has largely been overlooked by the photonics community. Recently, the ME has been predicted to arise in the diffusion process of light waves in certain photonic lattices under incoherent dynamics S18 . Such previous study indicated that, rather surprisingly, tightly-localized light waves can diffuse and reach uniform distribution in the lattice faster than weakly-confined fields, thus providing the first example of ME involving light waves. However, this effect concerning light diffusion is purely classical, prompting the key and intriguing question: do Mpemba-like effects for quantum light exist, i.e. originating from the quantum nature of light, which is fundamental to modern integrated photonic quantum systems?
In this Letter we predict the emergence of the Mpemba effect in quantum optics context originating from the quantum nature of light book , which does not have any classical counterpart. Specifically, we consider the decay dynamics of photon fields from a leaky optical resonator or waveguide R1 ; R2 ; R3 ; R3b ; R4 ; R5 ; Haus1 ; R6 ; R7 ; R8 (Fig.1), and show that the relaxation dynamics is strongly affected by the photon statistics of initially trapped light field. In particular, it is shown that bosonic ME arising from the quantum nature of light is observed when comparing the decay dynamics of coherent states (classical states of light) with certain non-classical states of light, such as Fock states, squeezed states and Schrödinger cat states.
@@
At the classical level the initial state is determined solely by the initial electric field amplitude $A(t=0)$ of the resonator mode, and the decay dynamics is the same regardless of the initial field amplitude. To unveil the occurrence of ME in the photon decay dynamics, we should consider the quantum nature of light. To this aim, let us assume that at initial time the photon field is described by the factorized state $|\psi(t=0)\rangle=|\psi_{S}(0)\rangle_{S}\otimes|0\rangle_{B}$, where $|0\rangle_{B}$ is the vacuum state of the zero-temperature reservoir (radiation) modes and
@@
and $A=A(t)$ is the classical decay law, i.e. the solution to Eq.(4) with the initial condition $A(t=0)=1$. The derivation of Eq.(9) is given in the Appendix A. Since $A(t)\rightarrow 0$ as $t\rightarrow\infty$, for any arbitrary initial probability amplitudes $\{p_{n}\}$ one has $\lim_{t\rightarrow\infty}\hat{\rho}_{S}(t)=|0\rangle_{S}\langle 0|\equiv\hat{\rho}_{S}^{\infty}$, i.e. as expected the asymptotic reduced density operator converges toward the vacuum state. The condition $A(t)\neq 0$ at any finite time $t$ ensures that in the transient relaxation process toward the stationary state it is not possible for the system to transiently cross the stationary state, which would introduce some ambiguities in the definition of the ME. The main point is that the relaxation dynamics strongly depends on the initial probability amplitudes $\{p_{n}\}$ of the photon field, which could result in the ME under suitable choice of initial states. To quantify how much the quantum state $\hat{\rho}_{S}(t)$ is far from $\hat{\rho}_{S}^{\infty}$ during the relaxation process, a distance measure $D(t)$ for density operators in Hilbert space has to be defined. Such a definition is not unique and ambiguities in the onset of ME could arise S5 ; S6c , however it is generally argued that any distance measure satisfying some minimal properties should result in a unique definition of the ME S5 . For quantum systems, the trace distance $D_{T}(t)$ and the Hilbert-Schmidt distance $D_{HS}(t)$ are generally used as possible metrics of distances in Hilbert space S9 ; S9b ; N1 ; N2 . They are defined by the relations
Figure 2: Bosonic Mpemba effect in photon decay dynamics with non-classical light. The figure compares the relaxation dynamics in the resonator/waveguide for an initial single-photon Fock state $|1\rangle_{S}=\hat{a}^{{\dagger}}|0\rangle_{S}$ (state I) and for the coherent state $|\alpha_{0}^{(coh)}\rangle_{S}$ (state II) with amplitude $\alpha_{0}=1$. Panels (a,b) show the behaviors of trace distance ($D_{T}$) and Hilbert-Schmidt distance ($D_{HS}$) versus normalized time $t/\tau_{c}$. A classical exponential decay law for $|A(t)|^{2}$ has been assumed, with photon lifetime $\tau_{c}$. The crossing between the two curves I and II, which arises at about $t/\tau_{c}\simeq 0.34$ for both metrics, is the signature of the ME. Panel (c) shows a few snapshots of the reduced density matrix elements $|\rho^{(S)}_{n,m}(t)|$ at some increasing times [from left to right: $t/\tau_{c}=0$ (initial states), $t/\tau_{c}=0.6$, $t/\tau_{c}=1.6$ and $t/\tau_{c}=3$]. Panels in the upper row refer to state I (Fock state decay), whereas panels in the lower row refer to state II (coherent state decay).
@@
Figure 3: Same as Fig.2, except that state I is the Schrödinger cat state $|\psi_{S}(t=0)\rangle_{S}=\mathcal{N}(|\alpha_{0}^{(coh)}\rangle_{S}+|-\alpha_{0}^{(coh)}\rangle_{S})$ with $\alpha_{0}=1.5$ ($\mathcal{N}$ is a normalization constant). Panels (a), (b) and (c) display the temporal behaviors of trace distance ($D_{T}$), Hilbert-Schmidt distance ($D_{HS}$), and the reduced density matrix elements $|\rho^{(S)}_{n,m}(t)|$, respectively. As compared to Fig.2, the crossing time of the two curves I and II depends on the metrics ($t/\tau_{c}\simeq 1.7$ in (a) and $t/\tau_{c}\simeq 1$ in (b)).
Figure 4: Same as Fig.2, except that state II is the squeezed vacuum state $|\psi_{S}(t=0)\rangle_{S}=\exp[(1/2)(\xi^{*}\hat{a}^{2}-\xi\hat{a}^{{\dagger}2})]|0\rangle_{S}$ with squeezing parameter $\xi=1$. Panels (a), (b) and (c) display the temporal behaviors of trace distance ($D_{T}$), Hilbert-Schmidt distance ($D_{HS}$), and the reduced density matrix elements $|\rho^{(S)}_{n,m}(t)|$, respectively.
@@
In conclusion, we predicted the existence of a bosonic Mpemba effect for light by considering photon decay dynamics of non-classical states of light. The analysis shows that certain non-classical states of light, such as photon number (Fock) states, decay faster than classical (coherent) light states when the full degrees of freedom of the photon field (i.e. reduced density matrix) are monitored. Our results unveil a type of photonic Mpemba effect of purely quantum nature, extending such a counterintuitive and curious phenomenon to the quantum optics context with potential interest in advanced quantum photonic technologies, such as quantum computing, quantum simulations, quantum metrology and quantum communications fuffa ; S17b , where fast relaxation dynamics could be desirable.